Beam Inc. bonds are trading today for a price of ​$1,299.78. The bond currently has 23 years until maturity and has a yield to maturity of 6.87​%. The bond pays annual coupons and the next coupon is due in one year. What is the coupon rate of the​ bond? (Round to one decimal​ place.)
Answer #1

Using financial calculator to calculate the coupon rate

Inputs: N= 23

I/y= 6.87%

Pv= -1,299.78

Fv= 1,000

Pmt= compute

We get, pmt as $95

Coupon rate = coupon / par value × 100

= 95 / 1,000 × 100

= 95 / 10

= 9.5%